OpenStudy (anonymous):

ACT MATH REVIEW QUESTION PLEASE HELP : a triangle with a perimeter of 66 inches has one side that is 16 inches long. The lengths of the other two sides have a ratio of 2:3. What is the length, in inches, of the longest side of the triangle?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

well, the perimeter is 66 inches and one side is 16 inches. our first step is to find the combined length of the other two sides first step: 66-16 = ?

OpenStudy (anonymous):

@kat0kitty

OpenStudy (anonymous):

66-16 =

OpenStudy (anonymous):

50

Vocaloid (vocaloid):

right, so we know that the other two sides have a combined length of 50 the 2:3 ratio tells us that one side is 2x and the other is 3x 2x + 3x = 50 solve for x

OpenStudy (anonymous):

x=10?

Vocaloid (vocaloid):

right, now we can plug x = 10 back into our other equations 2x = ? 3x = ?

OpenStudy (anonymous):

20 and 30

Vocaloid (vocaloid):

right. so, all this tells us that the three sides of the triangle are 16, 20, and 30 so the longest side is 30

OpenStudy (anonymous):

ok, thank you! would you mind helping me with another problem on ACT review? @Vocaloid

Vocaloid (vocaloid):

sure, post a new question and tag me

OpenStudy (anonymous):

ok, will do:-)